The value of the shunt resistance required to reduce the current passing through a galvanometer of resostance 48ohms to 1/5th of the original value is

Grade:12

1 Answers

noogler
8 years ago
S=G/n-1
n=i/i.......
given ig=i/5....n=5...............S=48/4=12ohms
so value ofshunt =12
approve if it helped u